### v3

# Design of supercurrent diode by vortex phase texture

Abstract
We investigate supercurrent nonreciprocal effects in a superconducting weak-link hosting distinct types of vortices. We demonstrate how the winding number of the vortex, its spatial configuration, and the shape of the superconducting lead can steer the sign and amplitude of the supercurrent rectification. We find a general criterion for the vortex pattern to maximize the rectification amplitude of the supercurrent. The underlying strategy is the search for specific vortex core positions yielding a vanishing amplitude of the supercurrent first harmonic. We also prove that supercurrent nonreciprocal effects can be used to diagnose high-winding vortex and to distinguish between different types of vorticity. Our results thus provide a toolkit to control the supercurrent rectification by means of vortex phase textures and nonreciprocal signatures to detect vortex states with nonstandard phase patterns.
I Introduction
Vortices represent fundamental topological excitations in superfluid and superconductors. They have been predicted and successfully observed in a broad range of systems including superconductors [1], liquid helium [2, 3, 4], ultracold atomic gases [5], photon fields [6], and exciton-polariton condensates [7, 8]. Vortices are generally characterized by a quantized phase-winding and a suppressed order parameter at their core. A gradient of the phase $\phi$ of the superconducting order parameter $\Delta=|\Delta|\exp{(i\phi)}$ yields a circulating supercurrent around the core of the vortex whereas the amplitude is vanishing, i.e. $|\Delta|\rightarrow 0$. For conventional superconductors, due to the single-valuedness of the superconducting order parameter, the winding number associated with the phase gradient is forced to be an integer $V_{0}$. In principle, $V_{0}$ can assume values larger than one. Apart from having $V_{0}$ vortices with winding number equal to one, a giant vortex with winding number $V_{0}$ can also realize. A giant vortex is expected to be relevant in small superconductors with confined geometries. To date, probing of such unconventional vortex state has been mostly addressed by magnetic means in suitably tailored geometric configurations, e.g. by Hall probe [9, 10, 11] and scanning SQUID microscopy [12], or by tunneling microscopy and spectroscopy [13, 14, 15, 16, 17].
Here, we unveil a specific relation between the occurrence of vortex states with any given winding number in a Josephson weak-link and the rectification of the supercurrent flowing across the junction. Supercurrent rectification is a timely problem at the center of intense investigation [18, 19, 20, 21, 22, 23, 24]. A large body of work devoted to supercurrent rectification focuses on superconducting states marked by linear phase gradients with, for instance, Cooper pair momentum [25, 26, 27], spin-flipper by ferromagnets [28], or helical phases [29, 30, 31, 32, 33], as well as screening currents [34, 35], and supercurrent related to self-field [36, 37] or back-action mechanisms [38]. Vortices or circular phase gradients, associated to conventional winding, are also expected to yield supercurrent diode effects, due to the induced Josephson phase shift [39], and their role has been investigated for a variety of physical configurations [37, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. However, whether and how superconducting phase patterns with nontrivial winding for the vorticity can be probed by nonreciprocal response, are problems not yet fully uncovered.
By studying nonreciprocal supercurrent effects arising from distinct types of vortex phase texture in Josephson weak-links, in this paper, we show that while a superconducting phase with vortices can generally lead to nonvanishing supercurrent rectification, the sign and amplitude of the rectification in the Josephson diode effect can be manipulated by the position of the vortex core and the winding of the phase vortex. We thereby uncover a general criterion to single out which phase vortex configuration can maximize the rectification amplitude of the supercurrent. Our findings provide a toolkit for the design and control of supercurrent rectification by vortex phase texture.
The structure of the paper is as follows. Sect. II provides the details of the model Hamiltonian, the formulation and the methodology employed in this study. In Sect. III we present the results for the supercurrent rectification in the case of one and two vortices within the superconducting leads for different values of the vortex winding. Sect. III is devoted to the analysis of the supercurrent rectification assuming that the supercurrent is vanishing at the boundary of the junction. In Section IV, we examine how the boundary affects the vortex phase texture and assess the resulting behavior of supercurrent rectification. The conclusions are reported in Sect. V.
II Model and methodology
In this section, we present the model Hamiltonian in real space for the superconducting leads, along with the methodology we used to analyze the current-phase relationship and to identify the maximum supercurrent that can flow in both directions across the junction.
Figure 1: (a) Schematic of a Josephson junction in the lattice model with a vortex phase texture. (b) A representative nonreciprocal current phase relation. (c)(d) Sketch of the superconducting weak link with a vortex placed on the left side of the junction. We show two representative vortex positions along the lateral direction. Since the supercurrent pattern is spatially modified by the phase vortex, it can become nonreciprocal, i.e. the forward supercurrent $I_{+}$ is different from the backward one $I_{-}$. The vortex winding, $V^{\mathrm{L}}_{0}$, can take any integer number $l$.
II.1 Model Hamiltonian
We consider a planar superconducting weak-link with the geometry shown in Fig. 1 (a). For convenience, the system size is expressed as $N_{y}a\times(N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x})a$ with $N_{y}$ being the lateral number of sites, $N^{\mathrm{L,R,N}}_{x}$ the site numbers in the left ($L$), right ($R$) and normal ($N$) region of the superconducting weak-link, and $a$ is the lattice length, respectively. The superconducting order parameter on the left and right side of the junction is given by $\Delta_{\mathrm{L,R}}e^{i\varphi_{\mathrm{L,R}}}$.
The Hamiltonian describing the superconducting junction illustrated in Fig. 1 (a) is written as
^$\hat{\mathcal{H}}=\sum_{j_{x},j_{y}}\sum_{j^{\prime}_{x},j^{\prime}_{y}}\hat{C}^{\dagger}(j_{x},j_{y})\hat{H}(j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y};\varphi)\hat{C}(j^{\prime}_{x},j^{\prime}_{y}),$ (1)
with $\varphi=\varphi_{\mathrm{L}}-\varphi_{\mathrm{R}}$ being the phase difference between the superconducting order parameters in the two sides of the junction and $\bm{j}=(j_{x},j_{y})$ with $j_{x}\in[-N^{\mathrm{L}}_{x}+1,N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x}]$ and $j_{y}\in[-(N_{y}-1)/2,(N_{y}-1)/2]$ indicating the site indices in the real space. Here, $\langle j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y}\rangle$ indicates the summation within the nearest-neighbor hopping and $\hat{C}^{\dagger}(j_{x},j_{y})=[c^{\dagger}_{j_{x},j_{y},\uparrow},c_{j_{x},j_{y},\downarrow}]$ denotes the creation operator at $\bm{j}$. Because we do not consider any symmetry breaking in the normal state, the number of the basis can be $2(N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+L^{\mathrm{R}}_{x})N_{y}$, not $4(N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+L^{\mathrm{R}}_{x})N_{y}$. Then $\hat{H}(j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y};\varphi)$ is given by
$\hat{H}(j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y};\varphi)$ $=\hat{H}^{\mathrm{L}}+\hat{H}^{\mathrm{R}}+\hat{H}^{\mathrm{N}}+\hat{H}^{\mathrm{L}}_{J}+\hat{H}^{\mathrm{R}}_{J},$ (2)
with the Hamiltonian in the Nambu space $\hat{H}^{\mathrm{L,N,R}}$ and the tunneling part $\hat{H}^{\mathrm{L,R}}_{J}$. The detail of the Hamiltonian is described by
$\hat{H}^{\mathrm{L}}$ $=\sum^{0}_{j_{x}=-N^{\mathrm{N}}_{x}+1}\sum^{(N_{y}-1)/2}_{j_{y}=-(N_{y}-1)/2}[(-\varepsilon)c^{\dagger}_{j_{x},j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}$
$-tc^{\dagger}_{j_{x},j_{y}+1,\uparrow}c_{j_{x},j_{y},\uparrow}-tc^{\dagger}_{j_{x}+1,j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}+$
$+\Delta_{\mathrm{L}}(j_{x},j_{y})c_{j_{x},j_{y},\uparrow}c_{j_{x},j_{y},\downarrow}]+h.c.,$ (3)
$\hat{H}^{\mathrm{N}}$ $=\sum^{N^{\mathrm{N}}_{x}}_{j_{x}=1}\sum^{(N_{y}-1)/2}_{j_{y}=-(N_{y}-1)/2}[(-\varepsilon)c^{\dagger}_{j_{x},j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}$
$-tc^{\dagger}_{j_{x},j_{y}+1,\uparrow}c_{j_{x},j_{y},\uparrow}-tc^{\dagger}_{j_{x}+1,j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}]+h.c.,$ (4)
$\hat{H}^{\mathrm{R}}$ $=\sum^{N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x}}_{j_{x}=-N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+1}\sum^{(N_{y}-1)/2}_{j_{y}=-(N_{y}-1)/2}[(-\varepsilon)c^{\dagger}_{j_{x},j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}$
$-tc^{\dagger}_{j_{x},j_{y}+1,\uparrow}c_{j_{x},j_{y},\uparrow}-tc^{\dagger}_{j_{x}+1,j_{y},\uparrow}c_{j_{x},j_{y},\uparrow}+$
$+\Delta_{\mathrm{R}}(j_{x},j_{y};\varphi)c_{j_{x},j_{y},\uparrow}c_{j_{x},j_{y},\downarrow}]+h.c.,$ (5)
with the on-site energy term $\varepsilon=-0.25t$, the energy gap function with spin-singlet $s$-wave state $\Delta_{\mathrm{L,R}}e^{\varphi_{\mathrm{L,R}}}$, and the nearest neighbor hopping terms $t_{x,y}$. In the present study, the pairing amplitude is on-site and included in the local term of Hamiltonian for the left and right-side superconductors. The tunneling Hamiltonian is expressed as
$\hat{H}^{\mathrm{L}}_{J}=\sum^{(N_{y}-1)/2}_{j_{y}=-(N_{y}-1)/2}[-t_{\mathrm{int}}tc^{\dagger}_{1,j_{y},\uparrow}c_{0,j_{y},\uparrow}]+h.c.,$ (6)
$\hat{H}^{\mathrm{R}}_{J}$ $=\sum^{(N_{y}-1)/2}_{j_{y}=-(N_{y}-1)/2}[-t_{\mathrm{int}}tc^{\dagger}_{-L^{\mathrm{L}}_{x}+L^{\mathrm{N}}_{x}+1,j_{y},\uparrow}c_{-L^{\mathrm{L}}_{x}+L^{\mathrm{N}}_{x},j_{y},\uparrow}]$
$+h.c.,$ (7)
with $t_{\mathrm{int}}$ the charge transfer amplitude at the interface setting out the transparency of the junction.
II.2 Josephson current and rectification
Next, we consider how the Josephson current flowing in the junction [Fig. 1 (a)] is evaluated. The current phase relation of the Josephson current is obtained by evaluating the variation of the free energy $F$ with respect to the phase bias across the junction:
$I(\varphi)=2\times\frac{2e}{\hbar}\frac{\partial F(\varphi)}{\partial\varphi},$ (8)
with the free energy in a Josephson junction at zero temperature evaluated as
$F(\varphi)$ $=\frac{1}{N_{x}N_{y}}\sum_{E<0}E(\varphi),$ (9)
$\hat{H}(j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y};\varphi)|\Phi\rangle$ $=E(\varphi)|\Phi\rangle.$ (10)
Here, $N_{x}=N^{\mathrm{L}}_{x}+N^{\mathrm{N}}_{x}+N^{\mathrm{R}}_{x}$ and $N_{y}$ denote the total number of sites along the $x$-direction and along the $y$-direction, and $E(\varphi)$ and $|\Phi\rangle$ stand for the eigenvalue and eigenstate of $\hat{H}(x,y;x^{\prime},y^{\prime};\varphi)$. We numerically obtain $E(\varphi)$ by the full diagonalization of $\hat{H}(j_{x},j_{y};j^{\prime}_{x},j^{\prime}_{y};\varphi)$ in Eq. (2). The computational analysis is performed at zero temperature, but a thermal change does not alter qualitatively the results. For our purposes, we recall that if time-reversal is broken then the supercurrent can be expanded in even and odd harmonics with respect to the phase bias variable as
$I(\varphi)=\sum_{m}[I_{m}\sin(m\varphi)+J_{m}\cos(m\varphi)],$ (11)
in Ref. [50]. To assess the nonreciprocal supercurrent due to the presence of the vortex in the junction, we evaluate the rectification amplitude $\eta$ that is conventionally expressed as
$\eta=\frac{I_{+}-|I_{-}|}{I_{+}+|I_{-}|},$ (12)
with $I_{+(-)}$ the maximum amplitude of the supercurrent for forward (backward) directions, respectively. For the presented results, we set the energy gap amplitude as $|\Delta_{0}|=0.02t$ ($t$ is the electron hopping amplitude), the maximum Josephson current value, without vortex, as $I_{\mathrm{0}}=0.012|\Delta_{0}|$, and the size of the superconducting leads as $N^{\mathrm{L}}_{x}=N^{\mathrm{R}}_{x}=\alpha N_{y}$ with $\alpha$ being the aspect ratio setting out a square or rectangular shape of the superconductors in the junction. The computation is performed for $N_{y}=30$, and $z^{\mathrm{L}}_{0}=10a$ for the size of the vortex core. A variation of these lengths does not affect the results, thus, for clarity, we introduce the coordinates $(j_{x},j_{y})$ in the Josephson junction [Fig. 2 (a)] (see details in Appendix A).
Figure 2: (a) Schematic illustration of the Josephson junction with real space coordinates. The analysis is performed for a vortex configuration having winding number $V^{\mathrm{L}}_{0}=+l$ with $l=1,2,3$. Evolution of the rectification amplitude $\eta$ with regard to the vortex core coordinates for different winding numbers: (b,c) $V^{\mathrm{L}}_{0}=1$, (f,g) $V^{\mathrm{L}}_{0}=2$, and (j,k) $V^{\mathrm{L}}_{0}=3$. Spatially resolved harmonics of the supercurrent: $I_{1}$ (red-solid line), $I_{2}$ (green-dotted), and $J_{1}$ (blue-dashed) indicate the odd-parity first harmonic, the odd-parity second harmonic, and the even-parity first harmonic amplitude, respectively. Longitudinal scan: $I_{1}$, $I_{2}$, and $J_{1}$ at a given $y^{L}_{0}$ as a function of $x^{L}_{0}$ for (d) $V^{\mathrm{L}}_{0}=1$, (h) $V^{\mathrm{L}}_{0}=2$, and (l) $V^{\mathrm{L}}_{0}=3$. Lateral scan: $I_{1}$, $I_{2}$, and $J_{1}$ at a given $x^{L}_{0}$ as a function of $y^{L}_{0}$ for (e) $V^{\mathrm{L}}_{0}=1$, (i) $V^{\mathrm{L}}_{0}=2$, and (m) $V^{\mathrm{L}}_{0}=3$. The gray dotted lines refer to the position of the maximal rectification and are a guide to indicate the values $I_{1}$, $I_{2}$, and $J_{1}$. We set $y^{\mathrm{L}}_{0}$ as (b)(d) $0.5a$ and (f)(h)(j)(l) $5.5a$, and $x^{\mathrm{L}}_{0}$ as (c)(e)(g)(i) $-23a$ and (k)(m) $-6a$. The aspect ratio is (b)-(i) $\alpha=3/2$ and (j)-(m) $\alpha=1$. The maximal rectification $\eta$ occurs for vortex core positions corresponding to a supercurrent with $I_{1}$, $I_{2}$, and $J_{1}$ components that are comparable in size. The sign change of $\eta$ is related to the vanishing of $J_{1}$ and to the zeros of $I_{1}$ when the amplitude is comparable to $J_{1}$. Multiple sign reversals of $\eta$ are observed for $V^{\mathrm{L}}_{0}=3$.
Figure 3: Amplitude of the first harmonics of the Josephson current versus the vortex core positions for different vortex winding: (a,b) $V^{\mathrm{L}}_{0}=1$, (c,d) $V^{\mathrm{L}}_{0}=2$, (e,f) $V^{\mathrm{L}}_{0}=3$. $\bar{I}_{1}$ and $\bar{J}_{1}$ indicate the odd- and even-parity first harmonics of the supercurrent evaluated by the direct Cooper pairs tunneling. Gray-dotted and dashed lines stand for the vortex core positions, $y^{\mathrm{L}}_{0}$ and $x^{\mathrm{L}}_{0}$, evaluated in Fig. 2. The aspect ratio is (a)-(d) $\alpha=3/2$ and (e,f) $\alpha=1$.
III Supercurrent Rectification: superconducting lead larger than the vortex phase coherence
In this section we study the supercurrent rectification by examining one and two vortices situated within the superconducting lead of the junction and assuming that it can be marked by different winding numbers. Here, we will first consider the case where the size of the vortex is much smaller than the superconducting lateral dimension so that the edges do not affect the vortex phase texture. Then, in Sect. IV we also analyze the rectification for the case of small size superconductors by imposing that the supercurrent is vanishing at the edge.
III.1 Single vortex configuration
To investigate the rectification properties of the supercurrent in the junction, we start by considering a single vortex configuration assuming a variable winding number and core position as shown in Figs. 1 (b,c). The vortex is placed on the left side of the junction but the results for vortices on the right side of the junction can be simply obtained using that inversion symmetry leads to a sign variation of the rectification amplitude (see Appendix A). Antivortex configurations through mirror and time-reversal transformations can also be directly deduced from the results of the single vortex (see Appendix A). To simulate the vortex state with the core at a given position, $\bm{r}^{\mathrm{L}}_{0}=(x_{0}^{\mathrm{L}},y_{0}^{\mathrm{L}})$, the phase value at the site $\bm{j}$ in the left-side superconductor is given by
$\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})=V^{\mathrm{L}}_{0}\arg[(j_{x}a-x_{0}^{\mathrm{L}}a)+i(j_{y}a-y_{0}^{\mathrm{L}})],$ (13)
with $V^{\mathrm{L}}_{0}$ being the winding number of the vortex. $\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ reverses its sign by changing the sign of $V^{\mathrm{L}}_{0}$, and $|V^{\mathrm{L}}_{0}|$ corresponds to the number of sign changes in the real space $\bm{j}$ for the phase value when winding around the core of the vortex. The pair potential $\Delta_{\mathrm{L}}$ and $\Delta_{\mathrm{R}}$ for spin-singlet $s$-wave state are given by
$\Delta_{\mathrm{L}}=|\Delta_{0}|\tilde{\Theta}(\bm{j},\bm{r}^{\mathrm{L}}_{0})e^{i\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})}$ (14)
and $\Delta_{\mathrm{R}}=|\Delta_{0}|$ with $|\Delta_{0}|$ being the superconducting energy gap amplitude. In the presence of a phase vortex texture, the amplitude of the pair potential $\Delta_{\mathrm{L}}$ is modified by $\tilde{\Theta}^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ for each site $\bm{j}$:
$\tilde{\Theta}^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})=\tanh\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right],$ (15)
with $z^{\mathrm{L}}_{0}=10a$ being the vortex size.
[…]
Moving to even high-winding, $V^{\mathrm{L}}_{0}=2$, we find that the maximal amplitude of the rectification occurs for vortex core positions that are now closer to the lateral edges of the superconductor but still midway from the interface [Figs. 2 (f,g)]. Inspection of the harmonics content of the current phase relation confirms that the sign change of the rectification amplitude $\eta$ occurs when $J_{1}$ reverses its sign [Figs. 2 (h,i)]. Furthermore, we find that the sign change of $\eta$ does not directly follow the $I_{1}$ sign change and the maximum of the rectification ($\eta\sim 25\%$) arises for vortex core positions whose $I_{1}$, $I_{2}$ and $J_{1}$ have comparable size. Let us then consider a vortex with odd high-winding number $V^{\mathrm{L}}_{0}=3$ assuming a square ($\alpha=1$) shape for the superconducting lead [Figs. 2 (j)-(m)]. For $V^{\mathrm{L}}_{0}=3$, we observe that the maximal values of the rectification amplitude occur for the vortex core position that is away from the mirror lines of the superconducting lead [Figs. 2 (j,k)]. The rectification can reach $30\%$ amplitude and there are multiple positions of the vortex core for which $\eta$ is vanishing [Figs. 2 (j,k)].
The analysis indicates that a key element to maximize the rectification is represented by the search of vortex configurations for which the first harmonics (even and odd-parity) are concomitantly almost vanishing. This is a general rule to get large rectification [51] (see also Appendix C). To this aim, one can make an analytical analysis of the first harmonics by considering the direct process of Cooper pairs transfer across the junction as given by [52, 50]
$\tilde{I}(\varphi,\bm{r}^{\mathrm{L}}_{0})$ $\propto\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\mathrm{Im}[\Delta_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})\Delta^{*}_{\mathrm{R}}].$ (16)
Taking into account the form of the superconducting order parameter we have that:
$\tilde{I}(\varphi,\bm{r}^{\mathrm{L}}_{0})$ $\propto\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tilde{\Theta}^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ (17)
×$\times[\sin\varphi\cos\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})+\cos\varphi\sin\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})].$
This relation is general and can be applied to any type of vortex phase texture. Here, from this expression, we can deduce how nonreciprocal supercurrent directly links to the winding of the phase vortex. Focusing on the coefficients of $\sin\varphi$ and $\cos\varphi$, the structure of the vortex indeed directly impacts on the even and odd-parity components of the first harmonics of the supercurrent. For convenience and in order to primarily extract the angular dependence of the harmonics content we neglect the amplitude variation in the vortex core. Hence, the effective first harmonics components $\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$ and $\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ of $\tilde{I}$ are substantially given by
$\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\cos\varphi_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (18)
$\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\sin\varphi_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ (19)
We checked that the spatial variation of the superconducting order parameter in the core of the vortex does not alter the qualitative conclusions of the analysis.
We find that their nodal lines can cross in different sites depending on the winding number and aspect ratio of the superconducting lead [Fig. 3]. In particular, for the case of even winding, $V^{\mathrm{L}}_{0}=2$, the breaking of $C_{4}$ rotational symmetry for a rectangular-shaped superconductor induces a shift of the nodal line of the odd-harmonic $\bar{I}_{1}$ towards the edge of the superconductor [Fig. 3 (c)]. Then, the crossing with the nodal lines for $\bar{J}_{1}$ shifts from $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-22a,0)$ to $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})\sim(-22a,\pm 10.5a)$. When considering a vortex with winding $V^{\mathrm{L}}_{0}=3$, due to the higher angular components, the crossings of the vanishing lines for $\bar{I}_{1}$ are in multiple points within the superconducting domain at $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})\sim(-7a,\pm 7.5a)$ and $(-22a,\pm 7.5a)$ [Fig. 3 (e,f)].
III.2 Two-vortex configuration
Figure 4: (a) Schematic illustration for two vortices in the left-side superconductor. (b,c) Rectification $\eta$ and (d,e) $I_{1}$, $I_{2}$, $J_{1}$ as a function of the literal direction $y^{\mathrm{L}}_{1}$ for (b,d) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$ and (c,e) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,-1)$. We set the core positions as $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-30,0.5)$ and $x^{\mathrm{L}}_{1}=-15$ ($N^{\mathrm{L}}_{x}=N^{\mathrm{R}}_{x}=45$, $N^{\mathrm{N}}_{x}=10$, $N_{y}=30$, $z^{\mathrm{L}}_{0}=z^{\mathrm{L}}_{1}=10a$). $I_{0}=0.122|\Delta_{0}|$ stands for the maximum Josephson current without any phase vortices in superconductors.
Next, we show the rectification caused by two vortices in one lead of the superconducting junction. The pair potential with two vortices in the left-side superconductor is expressed by:
$\Delta_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})=|\Delta_{0}|\prod^{1}_{l=0}\tilde{\Theta}_{l}(\bm{j},\bm{r}_{l})e^{i\varphi^{\mathrm{L}}_{\mathrm{v}l}(\bm{j},\bm{r}_{l})},$ (20)
with
$\tilde{\Theta}_{l}(\bm{j},\bm{r}_{l})=\tanh\left[\frac{|\bm{j}-\bm{r}^{\mathrm{L}}_{l}|}{z^{\mathrm{L}}_{l}}\right],$ (21)
the size of vortices $z^{\mathrm{L}}_{l}$, and the core positions $\bm{r}^{\mathrm{L}}_{l}=(x^{\mathrm{L}}_{l},y^{\mathrm{L}}_{l})$ for $l=0,1$. Each phase vortex is given by
$\varphi^{\mathrm{L}}_{\mathrm{v}l}(\bm{j})=V^{\mathrm{L}}_{l}\arg[(j_{x}a-x^{\mathrm{L}}_{l})+i(j_{y}a-y^{\mathrm{L}}_{l})],$ (22)
where $V^{\mathrm{l}}$ is the number of winding for each phase vortex.
We plot the rectification and Josephson components ($I_{1}$, $I_{2}$, and $J_{1}$) as a function of the literal direction $\tilde{y}^{\mathrm{L}}_{1}$ in a representative case [Fig 4]. Then we set the position of one phase vortex at $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-30a,0.5a)$. We choose each winding number as Fig. 4 (b,d) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$ and (c,e) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,-1)$. For $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$, the rectification is enhanced up to 40% near $y^{\mathrm{L}}_{1}=0$ [Fig. 4 (b)] owing to the comparable Josephson components $|I_{1}|\sim|I_{2}|\sim|J_{1}|$ [Fig. 4 (d)]. On the other hand, for $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,-1)$, because $|I_{1}|$ is larger than $|I_{2}|$ and $|J_{1}|$ [Fig. 4 (e)], the amplitude of the rectification is small compared with that for $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$ [Fig. 4 (c)]. Thus, the same sign for winding is favorable for enhancing the rectification.
Figure 5: (a,b) $\bar{I}_{1}$ and (c,d) $\bar{J}_{1}$ for the $\bm{j}^{\mathrm{L}}_{1}$ space at (a,c) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$ and (b,d) $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,-1)$. We set the core positions as $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-30a,0.5a)$ and $z^{\mathrm{L}}_{1}=10a$ ($N^{\mathrm{L}}_{x}=N^{\mathrm{R}}_{x}=45$ and $N_{y}=30$). The Gray line indicates the scanning site in Fig. 4.
Based on the discussion in the case of one phase vortex, we can also define $\bar{I}_{1}$ and $\bar{J}_{1}$ in this case. $\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})$ and $\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})$ are given by
$\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\cos\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1}),$ (23)
$\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\sin\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1}),$ (24)
with $\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})=\varphi^{\mathrm{L}}_{\mathrm{v}0}(\bm{j},\bm{r}^{\mathrm{L}}_{0})+\varphi^{\mathrm{L}}_{\mathrm{v}1}(\bm{j},\bm{r}^{\mathrm{L}}_{1})$. Fixing the position of one phase vortex at $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-30a,0.5a)$, we plot $\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})$ and $\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0},\bm{r}^{\mathrm{L}}_{1})$ for the $\bm{r}^{\mathrm{L}}_{1}$ space as shown in Fig. 5. For $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,+1)$ shown in Fig. 5 (a,c), since nodal lines appear around $x^{\mathrm{L}}_{1}\sim-11a$ for $\bar{I}_{1}$ and $y^{\mathrm{L}}_{1}\sim 0$ for $J_{1}$, the amplitude of $I_{1}$ and $J_{1}$ are small near these lines, respectively. For $(V^{\mathrm{L}}_{0},V^{\mathrm{L}}_{1})=(+1,-1)$, we obtain $\bar{I}_{1}\leq 0$ and the nodal line around $y^{\mathrm{L}}_{1}\sim 0$ for $J_{1}$ [Fig. 5 (b,d)]. Based on these $\bar{I}_{1}$ and $J_{1}$, $|I_{1}|$ is larger than $I_{1}$ and $J_{1}$ [Fig. 4 (d)]. Thus, using $\bar{I}_{1}$ and $\bar{J}_{1}$ is generic way to express $I_{1}$ and $J_{1}$.
IV Rectification with vanishing phase gradient at the edge
Figure 6: (a) Schematic illustration for the phase vortex in the superconducting lead with two image vortices to impose the vanishing of the supercurrent at the edge. Phase vortex with the boundary condition in a superconductor at (b) $V^{\mathrm{L}}_{0}=+1$, (c) $V^{\mathrm{L}}_{0}=+2$, and (d) $V^{\mathrm{L}}_{0}=+3$. The opposite circulation of the currents at the boundary is such that the phases have a vanishing gradient perpendicular to the edge. In panels (b), (c), (d) we have a vector field at a given point $(x,y)$ whose components are given by $[-\sin(\phi(x,y),\cos(\phi(x,y)]$, where $\phi(x,y)$ is the total phase due to the vortex and the image vortices shown in Eqs. (26), (29), and (30). The phase pattern consistently shows that there is no gradient when passing through the edges along the $y$ direction. For completeness, the pattern associated with the sum of the vortex vector fields is reported in Appendix E. Vortex core is set at $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-23a,-9.5a)$ in (b)-(d).
Figure 7: (a) Schematic illustration of the Josephson junction. At the edges (green lines) we assume that the component of the supercurrent perpendicular to the boundary is vanishing. The analysis is performed for a vortex configuration having winding number $V^{\mathrm{L}}_{0}=+l$ with $l=1,2,3$. Evolution of the rectification amplitude $\eta$ with regard to the vortex core coordinates for different winding numbers: (b,c) $V^{\mathrm{L}}_{0}=1$, (f,g) $V^{\mathrm{L}}_{0}=2$, and (j,k) $V^{\mathrm{L}}_{0}=3$. Spatially resolved harmonics of the supercurrent: $I_{1}$ (red-solid line), $I_{2}$ (green-dotted), $J_{1}$ (blue-dashed), and $J_{2}$ (cyan-dashed) indicate the odd-parity first harmonic, the odd-parity second harmonic, the even-parity first harmonic, and the even-parity second harmonic amplitude, respectively. Longitudinal scan: $I_{1}$, $I_{2}$, $J_{1}$, and $J_{2}$ at a given $j^{L}_{y0}$ as a function of $j^{L}_{x0}$ for (d) $V^{\mathrm{L}}_{0}=1$, (h) $V^{\mathrm{L}}_{0}=2$, and (l) $V^{\mathrm{L}}_{0}=3$. Lateral scan: $I_{1}$, $I_{2}$, $J_{1}$, and $J_{2}$ at a given $j^{L}_{x0}$ as a function of $j^{L}_{y0}$ for (e) $V^{\mathrm{L}}_{0}=1$, (i) $V^{\mathrm{L}}_{0}=2$, and (m) $V^{\mathrm{L}}_{0}=3$. The gray dotted lines refer to the site of the maximal rectification and are a guide to indicate the values $I_{1}$, $I_{2}$, $J_{1}$, and $J_{2}$. We set $y^{\mathrm{L}}_{0}$ as (b)(d)(j)(l) $0.5a$ and (f)(h) $9.5a$, and $x^{\mathrm{L}}_{0}$ as (c)(e)(k)(m) $-23a$ and (k)(m) $-28a$. The aspect ratio is $\alpha=3/2$. The maximal rectification $\eta$ occurs for vortex core positions corresponding to a supercurrent with $I_{1}$, $I_{2}$, and $J_{1}$ components that are comparable in size. The sign change of $\eta$ is related to the vanishing of $I_{1}$ and the zeros of $J_{1}$ when the amplitude is comparable to $I_{1}$.
In the previous section, we did not take into account the effects of the boundary near the edge on the vortex phase texture, as we focused on a scenario where the edge was sufficiently distant from the vortex core, making its impact on the phase texture negligible. However, this assumption does not hold if the vortex is positioned close to the edge or if the size of the superconductor is comparable to the superconducting coherence length. In such cases, the requirement for the supercurrent to be zero perpendicular to the edge can lead to a notable alteration of the superconducting phase surrounding the vortex core. One effective way to impose this constraint on the supercurrent amplitude at the edge of the junction is through the method of vortex images. As illustrated in Fig. 6 (a), when considering a single vortex in a superconductor, we can introduce two additional image vortices outside the superconductor, positioned at a specific distance from the edge to ensure that the supercurrent amplitude becomes vanishing at the edge. It is important to note that the image vortices must have the same winding as the vortex within the superconductor because using an image antivortex would not nullify the supercurrent; rather, it would increase the supercurrent at the edge, since both circulations would be moving in the same direction.
For taking into account the boundary condition for the supercurrent amplitude at the edge, we introduce the image vortices that lead to a modification of the phase value at each site in the superconducting leads as given by
$\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})\rightarrow\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}}),$ (25)
with
$\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})$ =$=\begin{cases}\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})+\varphi^{\mathrm{L}}_{a}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{a}})+(V^{\mathrm{L}}_{0}-1)\pi&j_{y}a>y^{\mathrm{L}}_{0}+(N_{y}-1)a/2\\ \varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})+(V^{\mathrm{L}}_{0}/2-1)\pi&y^{\mathrm{L}}_{0}-(N_{y}-1)a/2\leq j_{y}a\leq y^{\mathrm{L}}_{0}+(N_{y}-1)a/2\\ \varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{0}})+\varphi^{\mathrm{L}}_{b}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{b}})+V^{\mathrm{L}}_{0}\pi&j_{y}a<y^{\mathrm{L}}_{0}-(N_{y}-1)a/2\end{cases},$ (26)
$\varphi^{\mathrm{L}}_{a}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{a}})$ $=V^{\mathrm{L}}_{a}\arg[(j_{x}a-x_{a})+i(j_{y}a-y_{a})],$ (27)
$\varphi^{\mathrm{L}}_{b}(\bm{j},\bm{r}^{\mathrm{L}}_{\mathrm{b}})$ $=V^{\mathrm{L}}_{b}\arg[(j_{x}a-x_{b})+i(j_{y}a-y_{b})],$ (28)
where $V^{\mathrm{L}}_{a}=V^{\mathrm{L}}_{b}=V^{\mathrm{L}}_{0}$ are the winding number of the image vortices while the core positions $\bm{r}^{\mathrm{L}}_{a,b}$ are defined as
$x^{\mathrm{L}}_{a}$ $=x^{\mathrm{L}}_{0},$
$y^{\mathrm{L}}_{a}$ $=(N_{y}-1)a+N_{y}a-y^{\mathrm{L}}_{0}-2a,$
$x^{\mathrm{L}}_{b}$ $=x^{\mathrm{L}}_{0},$
$y^{\mathrm{L}}_{b}$ $=-y^{\mathrm{L}}_{0}+a.$
We introduce the modulation factors to obtain the continuous phase texture in a superconductor. In Fig. 6 (b)-(d), we plot the phase texture $\bm{F}=[f_{x}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),f_{y}(\bm{j},\bm{r}^{\mathrm{L}}_{0})]$ with
$f_{x}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ $=-\sin\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (29)
$f_{y}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ $=\cos\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (30)
in a superconductor at (b) $V^{\mathrm{L}}_{0}=+1$, (c) $+2$, and (d) $+3$. In the presence of two mirror vortices, we obtain $J_{y}(y=-14.5a)=J(y=14.5a)=0$ with
$J_{y}(y)$ $=f_{y}(y)-f_{y}(y\pm a),$ (31)
as the boundary condition.
Figure 8: Amplitude of the first harmonics of the Josephson current versus the vortex core positions for different vortex winding: (a,b) $V^{\mathrm{L}}_{0}=1$, (c,d) $V^{\mathrm{L}}_{0}=2$, (e,f) $V^{\mathrm{L}}_{0}=3$. $I_{1}$ and $J_{1}$ indicate the odd- and even-parity first harmonics of the supercurrent evaluated by the direct Cooper pairs tunneling. Gray-dotted and dashed lines stand for the vortex core positions, $y^{\mathrm{L}}_{0}$ and $x^{\mathrm{L}}_{0}$, evaluated in Fig. 5. The aspect ratio is $\alpha=3/2$.
By taking into account the modification of the superconducting phase due to the image vortices, in Figs. 7 (b)-(m), we present the rectification amplitude for different values of the vortex winding and by varying the vortex core position assuming $\alpha=3/2$ as representative cases for the aspect ratio of the superconducting lead. We also analyze the rectification with the boundary condition [Figs. 7 (b)-(m)] as we worked in Sect. III. For a conventional vortex with winding $V^{\mathrm{L}}_{0}=1$ in Figs. 2 (b,c), even with the boundary condition, the rectification has a similar tendency without the boundary condition [Figs. 2 (b,c)]. Other vortex core positions away from the symmetry lines also give a negligible rectification. To maximize the supercurrent rectification, the amplitude of the first harmonics in the current phase relation also plays a main role [Figs. 7 (d)(e)], as we discussed in Sect. III.
For the even high-winding, $V^{\mathrm{L}}_{0}=2$, we find that a small amplitude of the rectification occurs [Figs. 7 (f,g)] because of the large value of $I_{1}$ compared to $I_{2}$ and $J_{1}$ [Figs. 7 (h,i)]. For $V^{\mathrm{L}}_{0}=3$, we also observe that the maximal values of the rectification amplitude $\sim 30\%$ occur for the vortex core position that is near (away from) the mirror lines of the superconducting lead [Figs. 7 (j) (Figs. 7 (k))]. Their profile is qualitatively similar to those for $V^{\mathrm{L}}_{0}=1$ [Figs. 7 (d,e,j,k)], due to the boundary condition.
As discussed in Sect. III, we can also derive the analytical expression of the Josephson current:
$\tilde{I}(\varphi,\bm{r}^{\mathrm{L}}_{0})$ $\propto\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tilde{\Theta}^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ (32)
×$\times[\sin\varphi\cos\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})+\cos\varphi\sin\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})],$
by $\varphi^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})\rightarrow\tilde{\varphi}^{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$. Likewise, the effective first harmonics components $\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$ and $\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ of $\tilde{I}$ are given by
$\bar{I}_{1}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\cos\tilde{\varphi}_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (33)
$\bar{J}_{1}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum^{0}_{j_{x},j_{y}}\sin\tilde{\varphi}_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}).$ (34)
We find that their nodal lines can cross in different sites depending on the even and odd winding number [Fig. 8]. Due to the boundary condition, the same odd functions are qualitatively obtained for the odd winding number $V^{\mathrm{L}}_{0}=1,3$ [Figs. 8 (a,b,e,f)]. Even functions without (with) the nodal structure for $\bar{I}_{1}$ ($\bar{J}_{1}$) are demonstrated at $V^{\mathrm{L}}_{0}=2$ [Figs. 8 (c,d)].
V Conclusions
We have demonstrated that nonreciprocal supercurrents can be exploited to distinguish physical configurations with vortices having winding number equal to one, coexistence of vortices and antivortices or the occurrence of a giant vortex. A suitable platform to observe these effects can be based on Josephson junctions made of nanoislands in the presence of an applied magnetic field [13]. It is known that giant vortices can be induced by magnetic field for superconductors with coherence length much smaller than the magnetic penetration depth (e.g. Pb or Nb based nanostructures) [13]. Then, starting from a low magnetic field configuration with a vortex having, for instance, winding $V^{\mathrm{L}}_{0}=1$ nearby the center of the superconducting lead, the transition to a higher winding vortex state by the increase of the magnetic field will be accompanied to a sizable variation of the rectification. Detection of such transitions can be used to probe the high-winding vortex phase. Along this line, one can also envision a field-free platform [53] with circular current flow inducing high-winding vortices that concomitantly yield supercurrent rectification. Moreover, since a generic phase vortex texture can be expanded in harmonics by employing vortex configurations with different windings, then, in principle our results provide a toolkit to design the supercurrent rectification for a wide variety of superconducting phase patterns.
VI Acknowledgments
This work was funded by the EU’s Horizon 2020 Research and Innovation Framework Program under Grant Agreement No. 964398 (SUPERGATE), No. 101057977 (SPECTRUM), by the PNRR MUR project PE0000023-NQSTI, by the PRIN project 2022A8CJP3 (GAMESQUAD), and by the MAECI project “ULTRAQMAT”. We thank S. Ikegaya and Y. Tanaka for the valuable discussions.
Appendix A Phase vortex in real space
In Appendix A, we present the model and the methodology employed to calculate the Josephson current of the superconducting weak-link in the presence of vortices with different winding number.
The phase of the superconducting order parameter associated with the vortex configuration is represented in Fig. 9 (b). At the site $\bm{j}=(j_{x},j_{y})$, the phase vortex in the left-side superconductor is given by
$\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})=V^{\mathrm{L}}_{0}\arg[(j_{x}a-x^{\mathrm{L}}_{0})+i(j_{y}a-y^{\mathrm{L}}_{0})],$ (35)
with the position of the core $\bm{r}^{\mathrm{L}}_{0}=(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})$. In a similar way, one can introduce the phase vortex at the core $\bm{r}^{\mathrm{R}}_{0}=(x^{\mathrm{R}}_{x},y^{\mathrm{R}}_{0})$ in the right-side superconductor as
$\varphi^{\mathrm{R}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{R}}_{0})$ $=\pi-\bar{\varphi}^{\mathrm{R}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{R}}_{0}),$ (36)
$\bar{\varphi}^{\mathrm{R}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{R}}_{0})$ $=V^{\mathrm{R}}_{0}\arg[(j_{x}a-x^{\mathrm{R}}_{0})+i(j_{y}a-y^{\mathrm{R}}_{0})],$ (37)
due to the mirror symmetry in the $yz$-plane. Here, $V^{\mathrm{X}}_{0}$ denotes the vortex winding.
Figure 9: (a)(b) Phase of the vortex $\varphi^{\mathrm{X}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ and corresponding superconducting order parameter $\Theta^{\mathrm{X}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ in the real space with (a) X=L and (b) X=R. We also show the corresponding notations for the phase vortex for each side. $z^{\mathrm{X}}_{0}$ indicates the size of the vortex core wheareas the amplitude superconducting order parameter vanishes.
The pair potential is described as
$\Delta_{\mathrm{X}}=\begin{cases}\Theta(\bm{j},\bm{r}^{\mathrm{X}}_{0})e^{i\varphi^{\mathrm{X}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{X}}_{0})}&\text{with phase vortex}\\ |\Delta_{0}|&\text{without phase vortex}\end{cases},$ (38)
with the energy gap amplitude $|\Delta_{0}|$. In the former case, the energy gap amplitude $\Theta(\bm{j},\bm{r}^{\mathrm{X}}_{0}))$ is modified by the site:
$\Theta^{\mathrm{X}}(\bm{j},\bm{r}^{\mathrm{X}}_{0})$ $=|\Delta_{0}|\tilde{\Theta}^{\mathrm{X}}(\bm{j},\bm{r}^{\mathrm{X}}_{0}),$ (39)
$\tilde{\Theta}^{\mathrm{X}}(\bm{j},\bm{r}^{\mathrm{X}}_{0})$ $=\tanh\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{X}}_{0}|}{z^{\mathrm{X}}_{0}}\right],$ (40)
with the size of the vortex $z^{\mathrm{X}}_{0}$. We set the parameters as $\varepsilon=-0.25t$, $|\Delta_{0}|=0.02t$, and $t_{\mathrm{int}}=0.90$. In this study, we choose the size as $N^{\mathrm{N}}_{x}=10$, $N^{\mathrm{R}}_{x}=N^{\mathrm{L}}_{x}=\alpha N_{y}$, $N_{y}=30$ with the aspect ratio $\alpha$, and $z^{\mathrm{L}}_{0}=10a$.
In general, one can show that the following relations for the rectification amplitude $\eta$ holds when considering vortex (arbitrary winding amplitude $V_{0}>0$) or antivortex ($V_{0}<0$) configurations with a given position of the core $(x_{0},y_{0})$ and mirror related core positions:
$\eta(x_{0},y_{0},V_{0})=-\eta(-x_{0},y_{0},V_{0}),$ (41)
$\eta(x_{0},y_{0},V_{0})=-\eta(x_{0},-y_{0},V_{0}),$ (42)
$\eta(x_{0},y_{0},V_{0})=-\eta(x_{0},y_{0},-V_{0})\,.$ (43)
Therefore, regarding a combination of mirror and time-reversal symmetries, one can also obtain that a change from vortex to antivortex and of the vortex core position does not alter the sign and amplitude of the rectification:
$\eta(x_{0},y_{0},V_{0})=\eta(-x_{0},y_{0},-V_{0}).$ (44)
Appendix B Current phase relations in the presence of a single vortex
Figure 10: Current phase relation for (a,b) $V^{\mathrm{L}}_{0}=+1$, (c,d) $+2$, and (e,f) $+3$. We choose the position of the vortex core as (a) $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-23a,3.5a)$, (b) $(-20a,0.5a)$, (c) $(-23a,10.5a)$, (d) $(19a,10.5a)$, (e) $(-6a,10.5a)$, and (f) $(-34a,10.5a)$. Aspect ratio is (a)-(d) $\alpha=2/3$ and (e,f) $\alpha=1$. We also calculate the rectification: (a) $\eta=-0.07$, (b) $-0.20$, (c) $-0.07$, (d) $0.22$, (e) $-0.27$, and (f) $0.25$. Parameters: $\varepsilon=-0.25t$, $|\Delta_{0}|=0.02t$, $t_{\mathrm{int}}=0.90$, $N_{y}=30$, $z^{\mathrm{L}}_{0}=10a$, and $I_{0}=0.012|\Delta_{0}|$.
We show the representative cases of the current phase relation corresponding to the physical configurations Fig. 2 in the main text. Figure 10 presents the current phase relation for Fig. 10 (a)(b) $V^{\mathrm{L}}_{0}=1$, Fig. 10 (c)(d) $V^{\mathrm{L}}_{0}=2$, and Fig. 10 (e)(f) $V^{\mathrm{L}}_{0}=3$ at the position of a vortex core (a) $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-23a,3.5a)$, (b) $(-20a,0.5a)$, (c) $(-23a,10.5a)$, (d) $(19a,10.5a)$, (e) $(-6a,10.5a)$, and (f) $(-34a,10.5a)$. We obtain the nonreciprocal supercurrents caused by a phase vortex and these rectifications are Fig. 10 (a) $\eta=-0.07$, Fig. 10 (b) $-0.20$, Fig. 10 (c) $-0.07$, Fig. 10 (d) $0.22$, Fig. 10 (e) $-0.27$, and Fig. 10 (f) $0.25$, respectively.
Appendix C Condition to maximize the amplitude of the supercurrent rectification
Figure 11: Rectification $\eta$ of $I(\varphi)=I_{1}\sin\varphi+I_{2}\sin 2\varphi+J_{1}\cos\varphi+J_{2}\cos 2\varphi$ (a) for $(I_{1},J_{1})$ space and (b) as a function of $I_{2}$. We fix the parameters as (a) $(I_{2},J_{2})=(I_{0},0)$, (b) $(I_{1},J_{1},J_{2})=(I_{0},I_{0},0)$, (c) $(I_{1},J_{1})=(I_{0},I_{0})$, and (d) $(I_{1},J_{1},I_{2})=(I_{0},I_{0},-I_{0})$.
In Appendix C, we calculate the optimal conditions to maximize the supercurrent rectification starting from a generic current phase relation that includes the first harmonics (even and odd parity) and the second harmonic (odd parity). Let us then assume the following current phase relation:
$I(\varphi)=I_{1}\sin\varphi+I_{2}\sin 2\varphi+J_{1}\cos\varphi+J_{2}\cos 2\varphi.$
In order to find the optimal regimes for maximal rectification, one can evaluate $\eta$ as a function of the amplitudes $I_{1}$, $I_{2}$, and $J_{1}$. First, we calculate $\eta$ for $(I_{1},J_{1})$ at $(I_{2},J_{2})=(I_{0},0)$ in Fig. 11 (a). The maximum value of $|\eta|$ appears around $|I_{1}|=|J_{1}|\sim I_{0}$. Based on this point, we also calculate $\eta$ as a function of $I_{2}$ at $(I_{1},J_{1},J_{2})=(I_{0},I_{0},0)$ [Fig. 11 (b)]. The amplitude of $\eta$ is enhanced around $I_{2}=\pm 0.5I_{0}$. Likewise, to show the role of $J_{2}$, we calculate $\eta$ for $(I_{2},J_{2})$ at $(I_{1},J_{1}=(I_{0},I_{0})$ in Fig. 11 (c). $\eta$ is maximized around $J_{2}\sim 0$ for $|I_{2}|\sim I_{0}$. We also calculate $\eta$ for $J_{2}$ at $(I_{1},J_{1},I_{2})=(I_{0},I_{0},-I_{0})$ [Fig. 11 (d)]. The maximum $|\eta|$ appears at $I_{2}=0$. Thus, both $|I_{1}|$ and $|J_{1}|$ should be comparable with $|I_{2}|$ and $|J_{2}|$ should be minimized to maximize the rectification in Josephson current.
Appendix D Evaluation of first-harmonic component of supercurrent
We provide how we obtain the evaluation of the first-harmonics component of supercurrent in Appendix D. The first-harmonic components of the Josephson current at a site $\bm{j}$ can be estimated by considering the direct Cooper pair tunneling process as
$\tilde{J}(\varphi;\bm{j},\bm{r}^{\mathrm{L}}_{0})$ $\propto\mathrm{Im}[\Delta_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})\Delta^{*}_{\mathrm{R}}],$ (45)
where $\bm{r}^{\mathrm{L}}_{0}$ is the position of a vortex core in the left-side superconductor. By expanding the expression for $\tilde{J}(\phi)$, we obtain
$\tilde{J}(\varphi;\bm{j},\bm{r}^{\mathrm{L}}_{0})$
$\propto\mathrm{Im}\left[|\Delta_{0}|^{2}\tanh\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]e^{i(\varphi_{\mathrm{L}}-\varphi_{\mathrm{R}})}e^{i\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})}\right]$
$=|\Delta_{0}|\sin\varphi\tanh\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\cos\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$
$+|\Delta_{0}|\cos\varphi\tanh\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\sin\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (46)
with $\varphi=\varphi_{\mathrm{L}}-\varphi_{\mathrm{R}}$. The resulting first-harmonic Josephson current can be described by the spatial average of $\tilde{J}(\varphi;\bm{j},\bm{r}^{\mathrm{L}}_{0})$ over the number of sites in the superconducting leads:
$\tilde{I}(\varphi;\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tilde{J}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$ (47)
$\propto\sin\varphi\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tanh\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\cos\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$
$+\cos\varphi\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tanh\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\sin\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}).$
Based on this analysis, we get that
$\tilde{I}(\varphi;\bm{r}^{\mathrm{L}}_{0})$ $\propto F_{\mathrm{c}}(\bm{r}^{\mathrm{L}}_{0})\sin\varphi+F_{\mathrm{s}}(\bm{r}^{\mathrm{L}}_{0})\cos\varphi,$ (48)
with
$F_{\mathrm{c1}}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tanh\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\cos\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (49)
$F_{\mathrm{s1}}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tanh\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\sin\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}).$ (50)
Figure 12: Vector filed configuration from a phase vortex and two image vortices $\bm{g}=(g_{x},g_{y})$ for each site $\bm{j}=(j_{x},j_{y})$ for (a) $V^{\mathrm{L}}_{0}=1$, (b) $V^{\mathrm{L}}_{0}=2$, and (c) $V^{\mathrm{L}}_{0}=3$. Color bars indicate the norm of the vector. We choose the core position as $(x^{\mathrm{L}}_{0},y^{\mathrm{L}}_{0})=(-23a,-9.5a)$.
For completeness, we also note that the second harmonic Josephson current can be evaluated by considering the second-order Cooper pair tunneling processes:
$\tilde{I}_{2}(\varphi;\bm{r}^{\mathrm{L}}_{0})$
$\propto\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\mathrm{Im}[\Delta_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})\Delta^{*}_{\mathrm{R}}\Delta_{\mathrm{L}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})\Delta^{*}_{\mathrm{R}}]$ (51)
$=\sin 2\varphi\frac{|\Delta_{0}|^{4}}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tanh^{2}\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\cos 2\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})$
$+\cos 2\varphi\frac{|\Delta_{0}|^{4}}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tanh^{2}\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\sin 2\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}).$
Then, we can also find
$F_{\mathrm{c2}}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tanh^{2}\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\cos 2\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}),$ (52)
$F_{\mathrm{s2}}(\bm{r}^{\mathrm{L}}_{0})$ $=\frac{1}{N^{\mathrm{L}}_{x}N_{y}}\sum_{j_{x},j_{y}}\tanh^{2}\left[\frac{|\bm{j}a-\bm{r}^{\mathrm{L}}_{0}|}{z^{\mathrm{L}}_{0}}\right]\sin 2\tilde{\varphi}^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0}).$ (53)
Since both $I_{2}\propto F_{\mathrm{c2}}$ and $J_{2}\propto F_{\mathrm{s2}}$ are the $|\Delta_{0}|^{4}$ order, they are generally smaller in amplitude than $I_{1}$ and $J_{1}$, as they are proportional to $|\Delta_{0}|^{2}$. Then, close to the nodal lines of the first harmonics, the second harmonic components are comparable in amplitude.
Appendix E Phase pattern associated with the sum of the vortex and image vortex vector fields
Here, for completeness, we also show the phase pattern associated with the sum of the vortex and vortex-images vector fields as related to Fig. 7 (a). Based on the polar coordinates representation for the phase vortex, one can write that
$\bar{x}_{m}=\bar{r}_{m}\cos\varphi^{\mathrm{L}}_{m}(\bm{j},\bm{r}^{\mathrm{L}}_{m}),$ (54)
$\bar{y}_{m}=\bar{r}_{m}\sin\varphi^{\mathrm{L}}_{m}(\bm{j},\bm{r}^{\mathrm{L}}_{m}),$ (55)
with $\bar{x}_{m}=j_{x}a-x^{\mathrm{L}}_{m}$, $\bar{y}_{m}=j_{y}a-y^{\mathrm{L}}_{m}$, and $\bar{r}_{m}=\sqrt{\bar{x}_{m}^{2}+\bar{y}_{m}^{2}}$ for $m=\mathrm{v},a,b$. We consider the vector field described by three vectors $\bm{g}=(g_{x},g_{y})=(-\bar{y}_{\mathrm{v}}/\bar{r}_{\mathrm{v}}-\bar{y}_{a}/\bar{r}_{a}-\bar{y}_{b}/\bar{r}_{b},\bar{x}_{\mathrm{v}}/\bar{r}_{\mathrm{v}}+\bar{x}_{a}/\bar{r}_{a}+\bar{x}_{b}/\bar{r}_{b})$. Then this vector field $\bm{g}$ at the site $\bm{j}=(j_{x},j_{y})$ is given by
$g_{x}$ $=-\sin\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})-\sin\varphi^{\mathrm{L}}_{\mathrm{a}}(\bm{j},\bm{r}^{\mathrm{L}}_{a})-\sin\varphi^{\mathrm{L}}_{\mathrm{b}}(\bm{j},\bm{r}^{\mathrm{L}}_{b}),$ (56)
$g_{y}$ $=\cos\varphi^{\mathrm{L}}_{\mathrm{v}}(\bm{j},\bm{r}^{\mathrm{L}}_{0})+\cos\varphi^{\mathrm{L}}_{\mathrm{a}}(\bm{j},\bm{r}^{\mathrm{L}}_{a})+\cos\varphi^{\mathrm{L}}_{\mathrm{b}}(\bm{j},\bm{r}^{\mathrm{L}}_{b}).$ (57)
We plot the vector field $\bm{g}$ in real space at Fig. 12 (a) $V^{\mathrm{L}}_{0}=1$, Fig. 12 (b) $V^{\mathrm{L}}_{0}=2$, and Fig. 12 (c) $V^{\mathrm{L}}_{0}=3$, respectively.